Bob is 5 years older than Jenny. In 3 years he will be twice as old as her. What are their ages ?
Answer by jorel555(1290)

Let j be Jenny. Then Bob is j+5. So:
2(j+3)=j+5+3
2j+6=j+8
j=2
Jenny is 2; Bob is 7. ☺☺☺☺